OpenStudy (anonymous):

clarification plz write a system of equations. the sum of 2 numbers is 9. if one number is subtracted from the other, the result is 3 find the numbers

OpenStudy (anonymous):

so i got x+y=9 as equation 1

OpenStudy (anonymous):

is equation 2 x-y=3?

OpenStudy (anonymous):

if so the x=3+y

OpenStudy (anonymous):

so x = 9 and y =6?

OpenStudy (anonymous):

would i put it as 6, 9 or 9,6?

OpenStudy (anonymous):

@NeedMathHelpBadly any insight to this?

OpenStudy (anonymous):

I will help one sec:)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

hmm the answers cant be 9 and 6 becuase the problem is saying they add up to 9 and that is 15..... let me try to write an equation...

OpenStudy (anonymous):

(x + y ) = 9 + (x - y) = 3 ------------------ 2x = 12 x = 6 6 - y = 3 y = 3 now check 3 + 6 = 9 6 - 3 = 3

OpenStudy (anonymous):

Does this help? I got x=6 and y=3

OpenStudy (anonymous):

i see where i got 9 by mistake and helps alot

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem glad to help! Just message me if you have any more questions....